OpenStudy (anonymous):

integral(xsqrt(x-5)dx

OpenStudy (foolaroundmath):

\(x-5 = t^{2} \Rightarrow dx = 2tdt\) Try this substitution.

OpenStudy (anonymous):

idk how that would work i believe this is a parts integral

OpenStudy (zarkon):

u=x-5

OpenStudy (foolaroundmath):

\(x = 5+t^{2}\) Substituting this we get \(\int (5+t^{2})t.2tdt = \int (2t^{4} + 10t^{2})dt\) Can you take it off from here?

OpenStudy (anonymous):

yeah thanks man